OpenStudy (anonymous):

Use the Distributive Property to simplify: A.) (x+4)-2(x+5) B.) 1/2(4z+6) C.) (4+5)-(5+2) D.) x(x+7)

OpenStudy (anonymous):

simple just distribute the numbers or terms in front of the parentheses so for the first u would distribute -2 to all terms in the parentheses

OpenStudy (anonymous):

a*(b+c) = a*b + a*c does this help?

OpenStudy (anonymous):

for example we will take D) -> x(x+7) = x*x + 7x = x^2 + 7x

OpenStudy (anonymous):

try the other by yourself

OpenStudy (anonymous):

and let us know if you need more help :)